OpenStudy (anonymous):

find the domain and range of f(x) f(x)= sq. root 81-x^2

OpenStudy (anonymous):

domain: sqrt(u) is valid for u>= 0 so 81 - x^2 >=0 x^2 =< 81 -9=<x=<9 range: -(x^2) is always negative therefore 81 - x^2 is max at x = 0 min at x = +-9 0=<f(x)=<9